OpenStudy (anonymous):

(squaredrootof)x-5=(squaredrootof)x-1 [the -1 is not inside the squaredroot) :)

OpenStudy (dumbcow):

square both sides you have to FOIL the right side

OpenStudy (anonymous):

uhhh.... i need help with that :///

OpenStudy (anonymous):

\[\sqrt{x-5}=\sqrt{x}-1\]\[(\sqrt{x-5})^2=(\sqrt{x}-1)^2\]v-------------^\[(\sqrt{x}-1)(\sqrt{x}-1)\]FOIL \[(x-\sqrt{x}-\sqrt{x}+1)\]\[(x-2\sqrt{x}-1)\]\[x-5=x-2\sqrt{x}+1\]-x -1 -x -1 \[-6=-2\sqrt{x}\]/-2 /-2\[3=\sqrt{x}\]\[(3)^2=(\sqrt{x})^2\]\[9=x\]Finished!!!

OpenStudy (anonymous):

woah.. id nt get the foil.. why woudl (sqrdrt) of x multiplied by the same thing give x?

OpenStudy (anonymous):

I forget the name of the actual rule. But simply it's\[(x^2)(x^3)=x^5\]What you do is take the exponents and add them together.\[\sqrt{x}=x ^{1/2}\]Therefore\[(\sqrt{x})(\sqrt{x})=(x ^{1/2})(x ^{1/2})\]Which then makes\[(x ^{1/2})(x ^{1/2})=x^1\]or x.

OpenStudy (anonymous):

oh. haha. sorry :) and i don't get the x-1-x-1 part.. :(

OpenStudy (anonymous):

\[(x ^{a})(x ^{b})=x ^{a+b}\] That's the rule.

OpenStudy (anonymous):

I'm subtracting x and 1 from each side.

OpenStudy (anonymous):

ooohh... i see it now:) can you repeart that last part? after the little isue i just did? my computer won't show anything after that. :(

OpenStudy (anonymous):

after the -x-1-x-1?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

\[x-5=x-2\sqrt{x}+1\]-x -1 -x -1\[-6=-2\sqrt{x}\]/-2 /-2\[3=\sqrt{x}\]\[(3)^2=(\sqrt{x})^2\]\[9=x\]

OpenStudy (anonymous):

what do the /-2/-2 stand for?

OpenStudy (anonymous):

Divide by -2 on both sides.

OpenStudy (anonymous):

okay. :)

OpenStudy (anonymous):

I was taught "The Granny Rule: Granny's are fair, what they do to one chile they do to the other." Same with algebra, what you do to one side of the equation you have to do to the other.

OpenStudy (anonymous):

*child*

OpenStudy (anonymous):

mahaha:) that's a good way to remember it:D